OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of -6x^2 - 12x + 90 = 0. a.no real root b.two distinct real roots c.one distinct real root d.three distinct real roots

OpenStudy (anonymous):

The roots are x=-5 and +3. So answer is b.

OpenStudy (anonymous):

The reason is in this case D>0 (12^2-4(-6)(90) = 144+2160=2304>0